A firm recently paid a $0.30 annual dividend. The dividend is expected to increase by 10 percent in each of the next four years. In the fourth year, the stock price is expected to be $14.

If the required return for this stock is 13.5 percent, what is its current value? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

X Fax amount dividend at the end of year 1 (D1) = dividend at the end of year 2 (D2) = dividend at the end of year 3 (D3) = d

for formulas and calculations, refer to the image below -

1 x fc amount A с amount present value dividend at the end of year 1 (D1) = =0.3*1.1 =C3/1.135 dividend at the end of year 2

In case you have any query, kindly ask in comments.